OpenStudy (anonymous):

factor: m^3-m^2+2m-2?

OpenStudy (anonymous):

m^3 - m^2 + 2m - 2 = m(m^2 - 1) + 2(m - 1) = m(m + 1)(m - 1) + 2(m - 1) = (m - 1)[m(m+1)+2] = (m - 1)(m^2 + m + 2) Answer: (m - 1)(m^2 + m + 2)

OpenStudy (anonymous):

Whoa. That's not right. I have to fix it.

OpenStudy (anonymous):

m^3 - m^2 + 2m - 2 = m^2(m - 1) + 2(m - 1) = (m - 1)(m^2 + 2) Answer: (m - 1)(m^2 + 2)